A crane boom 10 m long, weighing 870 N and having its center of gravity 3m from the bottom, is hinged at the lower end and makes an angle of 30° with the vertical. It is held in position by a horizontal cable fastened at the upper end. The boom supports a 4350-N load at its upper end. Find the tension in the cable.


1
Expert's answer
2021-05-09T15:33:29-0400

"\text {let A lower point of the crane boom}"

"\text{B the upper point of the crane boom};F_b= 4350"

"Nb-\text{rope reaction force}"

"\text{C point of the center of gravity of the crane boom};F_c=870"

"\text{for hinge A}"

"\Sigma M_a(\vec{F_i})=0;"

"N_b*AB*\cos30\degree-F_b*AB*sin30\degree-F_c*AC*\sin30\degree=0"


"N_b= \frac{F_b*AB*sin30\degree+F_c*AC*\sin30\degree}{AB*\cos30\degree}="


"=\frac{4350*10*0.5 +870*3*0,5}{10*0.867}\approx 2662"


"\text{Aswer: } 2662\ N"